OpenStudy (anonymous):

How do I find the system of differential equations in p and u: P=a-T-betaU+hn n(dot)=1(P-n) U(dot)=-k(m-p)

OpenStudy (experimentx):

is it time dependent?

OpenStudy (anonymous):

yes n(dot) and U(dot) are. and I assume that the result should be time dependent too.

OpenStudy (experimentx):

and a, T and h are constants??

OpenStudy (anonymous):

yes. a,beta, T bigger than 0 and h lies between 1 and 0, k bigger than 0

OpenStudy (anonymous):

I am not sure where to start. should I just shift the equations first that I have u= and p= on the left hand side?

OpenStudy (experimentx):

this one right?? \[ P = a-T-\beta U+hn \]

OpenStudy (anonymous):

yes. So should I just take this equation and derive P with respect to time. but than we want a small p in the result. Do you think it has to be adjusted?

OpenStudy (anonymous):

in the first task we have to find the system of differential equation in p and u, and in the second task we need to draw a phase diagram, but this is less of a problem.

OpenStudy (experimentx):

what is it related to??

OpenStudy (anonymous):

I have difficulties to find the system of differential equations, but if I have it tho draw a phase diagram wont be difficult

OpenStudy (experimentx):

i meant the system? kinda looks like physics!!

OpenStudy (anonymous):

ouuu no it is just mathematics for economist. and the task is without any implications to an economical problem, that is why I m kinda stucked...and not sure where to start..

OpenStudy (experimentx):

\[ P = a-T-\beta U+hn \] \[ \dot{n}=1(P-n)\] \[ \dot{U}=-k(m-p) \] any relation between P and p??

OpenStudy (anonymous):

there is nothing about that in the task. So probably not

OpenStudy (anonymous):

and n is actually a pi

OpenStudy (anonymous):

so it could be that P stand for normal price, U-utility, and pi- profit

OpenStudy (anonymous):

so that p is real price...but there is nothig about in the task and just a guess

OpenStudy (experimentx):

\[ \dot{P} = -\beta \dot{U} + h \dot{\pi} = -k(m-p)+(P-\pi)\] there is still \( \pi \) ... i must be function of of P .. i.e \( \pi(P) \) if it is not then .. it might be incorrect. also is \( -k(m-p) \) all constants??

OpenStudy (anonymous):

no p is not a constant, we have to find the system of differential equation in p and u, so p and u must be dependent on time

OpenStudy (experimentx):

is it a partial differential equation??

OpenStudy (anonymous):

it must be total

OpenStudy (experimentx):

are you sure this does not involve partial derivative?? \[ \dot{\pi}=1(P-\pi)\] solving this equation, you get, \( \pi = C_1 e^{\int dt} + e^{\int +P dt} \int P e^{\int dt} dt\) You can substitute the value of \( \pi\) in above equation ... But I am not sure ... http://www.codecogs.com/latex/eqneditor.php Try using latex

OpenStudy (anonymous):

I m really not sure

OpenStudy (experimentx):

can you rewrite all equations?? in latex?? .. also make sure if one quantity is dependent on other variables or not??

OpenStudy (anonymous):

I am not familiar with latex. but the 3 equations you tiped in before where correct, and those three equations are all I have to solve this task. I wish I would have more hints.

OpenStudy (anonymous):

from these 3 equations I have to find the system of differential equations in p and u.

OpenStudy (experimentx):

Ops sorry ... no P there, \[ \pi = C_1 e^{\int dt} + e^{\int dt} \int P e^{\int dt} dt\] This is getting difficult .. P is again functio of pi

OpenStudy (anonymous):

hmm but what is the system of differential equation? and shouldt there be a solution for p and u?

OpenStudy (experimentx):

let's google it ... what is the name of this equation??

OpenStudy (experimentx):

Sorry ... i guess i can't help it!!

OpenStudy (anonymous):

I tried to google it and links to this side are shown. But thanks though, I ll try to solve it. maybe it works somehow ;)

OpenStudy (experimentx):

k ..